OpenStudy (anonymous):

Solve the equation. 8x2 – 4 = 28 A. 2, –2 B. 4, –4 C. 8, –8 D. 16, –16

OpenStudy (anonymous):

@Hero

hero (hero):

1. Add four to both sides 2. Divide both sides by 8 3. Square root both sides Don't forget to include the \(\pm\) in front of square root on the right sides.

OpenStudy (anonymous):

I lost you at 3.............

hero (hero):

You know what a square root is....\(\sqrt{\space\space\space}\) Put x^2 and 16 under it

hero (hero):

And don't forget the \(\pm\) symbol

hero (hero):

Let me post the steps. I don't want to confuse you

OpenStudy (anonymous):

Ok but what about the multiple choice it doesnt show an equation?

hero (hero):

\(8x^2 - 4 = 28\) Add four to both sides: \(8x^2 = 32\) Divide both sides by 8: \(x^2 = 16\) Square root both sides: \(x = \pm\sqrt{16}\) \(x = \pm 4\)

OpenStudy (anonymous):

Its b

OpenStudy (anonymous):

ok I get the formula now thanks @Hero

OpenStudy (anonymous):

Its A.